OpenStudy (anonymous):

the function f(x,y) log_2_(4+x) + log_2_(3y-1) + 1- log_2_(x) can be rewritten to the form F(x.y) Log_2_(p). Determine P

OpenStudy (loser66):

\[log_2 P= log_2 (4+x)+log_2 (3y-1)+1-log_2(x)\] right?

OpenStudy (anonymous):

yesss

OpenStudy (loser66):

just replace \(1= log_2 2\) then combine them by property of log and you get the answer

OpenStudy (loser66):

\[log_2 P= log_2 (4+x)+log_2 (3y-1)+log_2 2-log_2(x)\] and apply \(log_2 a + log_2 b= log_2 (a*b)\) and \(log_2 c - log_2 d) =log_2(\dfrac{c}{d})\)

OpenStudy (anonymous):

i have p= \[\frac{24y+8+6xy-2x }{ x }\]

OpenStudy (anonymous):

i'm not sure if i have to do something further